• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name lkaufman

    Log entry time 08:03:07 on September15,2005

    Entry number 154780

    keyword=Owl Shift Summary

    Owl Shift Summary
    -----------------------
    Summary: Smooth production running all morning.
    -----------------------
    00:00 - Production running.
    
    01:44 - IHWP flippered to in
    
    06:30 - Profile scan
    
    06:45 - IHWP flippered to out
    
    Compton runs 12180 and 12181 measure the beam polarization to be 84%.
    
    SUMMARY OF RUNS
    --------------------
    5060 - production, IHWP out, 44.8k good pairs
    5061 - production, IHWP out, 49.1k good pairs
    5062 - production, IHWP in, 46.4k good pairs
    5063 - production, IHWP in, 48.9k good pairs
    5064 - production, IHWP in, 47k good pairs
    5065 - production, IHWP in, 44.4k good pairs
    5066 - production, IHWP in, 46k good pairs
    5067 - profile scan
    5068 - production, IHWP out, 47.1k good pairs
    5069 - production, IHWP out, passed on to day shift